OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (hoblos):

parallel lines have same slope so the slope is m=2 use the formula y-y1=m(x-x1) y-3=2(x-5) y-3 = 2x -10 2x-y=7

OpenStudy (anonymous):

so the slope is 2 because its next to the x right? y=mx +b?

OpenStudy (hoblos):

yes you are right =)

OpenStudy (anonymous):

yes thanks :) & how would i do one if they asked me to show them a perpendicular one for example my next problem is this: Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (hoblos):

the product of slopes of two perpendicular lines is (-1)

OpenStudy (hoblos):

in this question the slope of the line would be -1/2

OpenStudy (anonymous):

so its y = mx + b, what would be b for perpendicular lines?

OpenStudy (anonymous):

like would b be 3 or 6?

OpenStudy (hoblos):

use the formula y-y1=m(x-x1) where x1 and y1 are the coordinates of a point on the line here x1=6 and y1=3

OpenStudy (anonymous):

i dont know how to use that formula :/

OpenStudy (hoblos):

y-3=-1/2(x-6)

OpenStudy (hoblos):

now just simplify

OpenStudy (anonymous):

now what do i do with that? i got y = 2 & 2x -12

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). a)y = 2x + 6 b)y = - x + 3 c)y = - x + 6 d)y = 2x + 3

OpenStudy (anonymous):

thats the problem but i know its either a or b i just dont know which

OpenStudy (hoblos):

in option (c) is it -x or -(1/2)x

OpenStudy (anonymous):

i copied it wrong my bad

OpenStudy (anonymous):

in option b its -1/2x +3

OpenStudy (anonymous):

& option c its -1/2x +6

OpenStudy (hoblos):

so if you simplify y-3=-1/2(x-6) you will get y= -1/2x +6 so it is c

OpenStudy (anonymous):

thankssss so much ! are you busy right now ?

OpenStudy (hoblos):

yw no im not...

OpenStudy (anonymous):

you have fb i can msg better on ? i have like a couple more doubts :) ? lol

OpenStudy (hoblos):

yes i have whats your name?? mine is ahmad hoblos =)

OpenStudy (anonymous):

if i found the one whos actually you i sent you a request right now :)